OpenStudy (anonymous):

Replace the ? with a polynomial to make a true statement: 2x -2 + x -1-3 = x -2 (?)

OpenStudy (anonymous):

? = 2x -4

OpenStudy (anonymous):

Solving for ?: 2? = x - (2x - 2 + x - 1 - 3) = x -(3x - 6) = 6 - 2x ? = 3 - x.

OpenStudy (anonymous):

2x + x - 2 - 1 - 3 = 3x - 6 3x - 6 = x - 2 + (2x - 4)